# LLM-Based Compact Reranking with Document Features for Scientific Retrieval

### v1

@@ Abstract
Scientific retrieval is essential for advancing academic discovery. Within this process, document reranking plays a critical role by refining first-stage retrieval results. However, standard LLM listwise reranking faces unique challenges in the scientific domain. First-stage retrieval is often suboptimal in the scientific domain, so relevant documents are ranked lower. Moreover, conventional listwise reranking inputs the full text of candidates into the context window, limiting the number of candidates that can be considered. As a result, many relevant documents are excluded before reranking, constraining overall retrieval performance. To address these challenges, we explore compact document representations based on semantic features (e.g., categories, sections and keywords) and propose CoRank, a training-free, model-agnostic reranking framework for scientific retrieval. It presents a three-stage solution: (i) offline extraction of document-level features, (ii) coarse reranking using these compact representations, and (iii) fine-grained reranking on full texts of the top candidates from (ii). This hybrid design provides a high-level abstraction of document semantics, expands candidate coverage, and retains critical details required for precise ranking. Experiments on LitSearch and CSFCube show that CoRank significantly improves reranking performance across different LLM backbones (nDCG@10 from $32.0$ to $39.7$). Overall, these results highlight the value of information extraction for reranking in scientific retrieval. Code will be publicly available.
LLM-Based Compact Reranking with Document Features for Scientific Retrieval
@@
Figure 1: Instead of directly reranking full documents, we extract features, rerank a larger candidate pool with them, and finally refine the shortlist with full documents.
1 Introduction
@@
Scientific retrieval (Lawrence et al., 1999; White et al., 2009) is crucial for scientific discovery. While current retrievers are effective at retrieving broadly relevant scientific papers, they often struggle to differentiate between documents covering similar topics (Sciavolino et al., 2021; Liu et al., 2021), making fine-grained relevance estimation essential. Therefore, the reranking stage (Carbonell and Goldstein-Stewart, 1998; Kurland and Lee, 2005) is particularly important (Gao et al., 2021b), as it refines the first-stage retrieval to better distinguish between closely ranked documents.
@@
Recently, large language models (LLMs) (Grattafiori et al., 2024; OpenAI et al., 2024) have significantly advanced document reranking, particularly through their application to listwise reranking (Sun et al., 2023a; Ma et al., 2023). In this setting, LLMs jointly assess a set of retrieved candidates within their context window and generate a reordered ranking based on their relevance. Previous works (Pradeep et al., 2023a; Gangi Reddy et al., 2024; Liu et al., 2024c; Ren et al., 2024) show that LLM-based listwise rerankers outperform prior embedding-based approaches (Nogueira et al., 2019b, 2020) on benchmarks like BEIR (Thakur et al., 2021).
@@
The standard practice in LLM listwise reranking (Sun et al., 2023a; Pradeep et al., 2023a, b; Gangi Reddy et al., 2024; Liu et al., 2024c) is to input the full text of each candidate document into the context window. However, this approach faces key limitations in scientific retrieval. In the scientific domain, retrievers often show limited performance (Kim et al., 2023; Kang et al., 2024a, b), so truly relevant documents may not rank high enough in the first-stage retrieval. Moreover, due to the substantial token overhead of full text representation and finite context length, rerankers can only operate on a limited number (typically 20 per prompt) (Sun et al., 2023a; Ma et al., 2023) of candidates. As a result, when first-stage retrieval is suboptimal, the reranking performance is inherently constrained (Reddy et al., 2023).
To address these limitations, we explore an alternative document representation that is both compact and informative. Instead of using full text, we investigate an information extraction (IE)–based approach (Niklaus et al., 2018; Zhou et al., 2022; Liu et al., 2022), representing each scientific paper using high-level features such as categories, sections, and keywords. We find that this representation significantly improves per-document token efficiency, enabling a larger pool of candidates to fit within the LLM’s context window. This, in turn, makes the reranking process more robust to suboptimal first-stage scientific retrieval results. It also simplifies the input by filtering out irrelevant details, making it easier for LLMs to interpret. However, since IE cannot guarantee complete coverage of all relevant content, full-text inputs still serve as a valuable complement when fine-grained relevance comparisons are required.
@@
Motivated by these insights, we propose CoRank, a training-free, model-agnostic reranking framework for science retrieval. CoRank consists of three stages: (i) Offline Preprocessing: extract high-level semantic features like categories and keywords from unstructured scientific documents; (ii) Coarse Reranking: use these compact representations for an initial ranking and select a subset of top candidates; (iii) Fine-grained Reranking: rerank the top candidates using full scientific documents to recover the potentially missing details during information extraction. Our design enhances the robustness against suboptimal first-stage retrieval with compact feature representation and maximizes the overall effectiveness with final full text reranking.
We evaluate CoRank on two scientific retrieval benchmarks: LitSearch (Ajith et al., 2024) and CSFCube (Mysore et al., 2021). Empirical experiments show that across various LLM backbones, CoRank achieves an absolute improvement of $+11.5$ nDCG@10 without the sliding window strategy, and retains a $+3.9$ gain with it. These results demonstrate significant effectiveness of our method for reranking in scientific retrieval. Overall, our contributions are threefold:
@@
    We are the first to reveal the unique limitations of LLM listwise reranking in scientific retrieval.
@@
    We are the first to explore semantic features as compact representations in reranking and propose CoRank, a framework based on this design.
@@
    We show that CoRank significantly improves reranking performance in scientific retrieval.
These findings underscore the value of compact feature extraction as a pre-analysis step for improving the reranking performance in scientific retrieval.
@@
Figure 2: Overview of our feature extraction pipeline: from unstructured documents, we apply zero-shot LLM information extraction to obtain document features including categories, sections, pseudo queries, and keywords, which are then combined into compact representations.
2 Preliminary Analysis
@@
In this section, we first define listwise reranking, then show the limitations of current full text-based methods and explore compact and informative document representation options for reranking in the scientific domain.
2.1 Problem Definition
Reranking Input.
@@ 2.1 Problem Definition
Given a query $q$ and a large corpus $\mathcal{P}_{n}$ of $n$ scientific papers, a first-stage retriever selects a ranked list of $m$ candidate documents $\mathcal{C}=\{p_{1},p_{2},\dots,p_{m}\}\subset\mathcal{P}_{n}$. Typically we have $m\ll n$ due to the scale of the corpus and the finite capacity of rerankers. This candidate list $\mathcal{C}$ is then passed to a reranking model, which aims to reorder the documents such that more relevant ones are ranked higher.
Reranking Objective.
@@
The goal of document reranking is to find a permutation $\sigma:\{1,\dots,m\}\!\to\!\{1,\dots,m\}$ whose application
@@
orders documents in more accurate descending relevance to $q$. The quality is often (Sun et al., 2023a; Gangi Reddy et al., 2024; Liu et al., 2024c) evaluated using top-$k$ metrics such as nDCG@10, with small $k$ values that stress accuracy at the top of the ranked list.
LLM Listwise Reranking.
Large language models can inspect $m$ candidate documents in their context window and generate a permutation that reflects their relevance ordering. Formally, the model acts as follows:
$\sigma_{\text{LLM}}\;=\;\mathrm{LLM}(q,\,\mathcal{C})\;\in\;S_{m},$
where $S_{m}$ is the set of all permutations on $\{1,\dots,m\}$. Applying this permutation to the candidate list yields
@@
2.2 Current Limitations
The standard approach in LLM-based listwise reranking (Sun et al., 2023a; Pradeep et al., 2023a, b; Gangi Reddy et al., 2024; Liu et al., 2024c) feeds the full text of each candidate into the model’s context window. While this allows LLMs to capture the complete content of each document, it incurs some limitations, especially in scientific retrieval.
Suboptimal First-Stage Retrieval.
Unlike in general-domain scenarios, first-stage retrievers—whether sparse (Robertson et al., 2009; Nogueira et al., 2019a) or dense (Izacard et al., 2021; Wang et al., 2022)—struggle to generalize to the scientific domain (Thakur et al., 2021; Bonifacio et al., 2022). This is due to the complexity of long-tail concepts (Kang et al., 2024b, 2025) and the lack of large-scale supervised training data (Bonifacio et al., 2022; Li et al., 2023). Consequently, their retrieval performance is limited, and truly relevant documents are often ranked much lower (Ajith et al., 2024).
Token Consumption.
Full text representation also introduces significant token overhead, with individual scientific papers often consuming hundreds or even thousands of tokens (Thakur et al., 2021; Ajith et al., 2024; Mysore et al., 2021). At the same time, LLMs have a limited effective context length (Dai et al., 2019; Xiong et al., 2024) and are known to suffer from issues such as positional bias for long inputs (Liu et al., 2024a; Tian et al., 2024). As a result, full-text rerankers are constrained to operate on a small number of candidates, typically 20 documents per input prompt (Sun et al., 2023a; Ma et al., 2023; Pradeep et al., 2023a, b).
@@ 2.2 Current Limitations
2.3 Semantic Features as Alternatives
Given the limitations of full-text representations in scientific-domain reranking, we explore an alternative approach, which is based on document-level semantic features such as categories and keywords. The intuition is that these features are both significantly more concise and capable of preserving the core semantics. Therefore, more candidate documents can be effectively considered during reranking. Specifically, we examine the following four types of document semantic features:
Category.
Categories (Sun et al., 2023b; Zhang et al., 2024) offer a high-level topical overview of scientific documents. Each document is assigned a three-level hierarchical category path in the format {Category} → {Subcategory} → {Subsubcategory}, which provides a broad-to-specific classification on its topic.
Section.
Sections (Zhou et al., 2023) consist of multiple subtitle-style strings, each summarizing a major part of the scientific document. They capture the internal structure of the document and serve as mid-level semantic signals that enhance category-level summaries.
Keyword.
Keywords (Rose et al., 2010; Lee et al., 2023) are terms or entities that represent fine-grained lexical concepts within a document. They provide the most specific information among different granularities.
Pseudo Query.
Pseudo queries (Sachan et al., 2022; Kang et al., 2025) are synthetic user questions based on the content. This feature offers a unique query-aligned perspective of the document, simulating how the document might be retrieved in real-world scenarios.
These features cover most common types of document-level IE and vary in granularity and style. All semantic features are extracted using LLM-based zero-shot information extraction. The prompt templates used for this extraction are provided in Appendix A.
Among these features, sections, keywords, and pseudo queries have multiple elements for each document. To reduce noise and focus on the most relevant content, we apply an adaptive selection strategy at inference time. We compute dense embedding similarities between the query and each extracted element and retain only the most relevant: 5 keywords (from 30), 1 pseudo query (from 20), and 1 section (from 3). This improves content relevance while minimizing token overhead. Ablation of this strategy can be found in Section 4.2.
Figure 3: Token efficiency and performance comparison across different document representations. (a) Per-document token lengths distribution. (b) nDCG@10 scores for different number of documents in the context window of single LLM input (c) Context token overhead to reach equal reranking performance.

### v2

# CoRank: LLM-Based Compact Reranking with Document Features for Scientific Retrieval

@@ Abstract.
Scientific retrieval is essential for advancing scientific knowledge discovery. Within this process, document reranking plays a critical role in refining first-stage retrieval results. However, standard LLM listwise reranking faces challenges in the scientific domain. First-stage retrieval is often suboptimal in the scientific domain, so relevant documents are ranked lower. Meanwhile, conventional listwise reranking places the full text of candidates into the context window, limiting the number of candidates that can be considered. As a result, many relevant documents are excluded before reranking, constraining overall retrieval performance. To address these challenges, we explore semantic-feature-based compact document representations (e.g., categories, sections, and keywords) and propose CoRank, a training-free, model-agnostic reranking framework for scientific retrieval. It presents a three-stage solution: (i) offline extraction of document features, (ii) coarse-grained reranking using these compact representations, and (iii) fine-grained reranking on full texts of the top candidates from (ii). This integrated process addresses suboptimal first-stage retrieval: Compact representations allow more documents to fit within the context window, improving candidate set coverage, while the final fine-grained ranking ensures a more accurate ordering. Experiments on 5 academic retrieval datasets show that CoRank significantly improves reranking performance across different LLM backbones (average nDCG@10 from $50.6$ to $55.5$). Overall, these results underscore the synergistic interaction between information extraction and information retrieval, demonstrating how structured semantic features can enhance reranking in the scientific domain.
Information Retrieval; Scientific Document Search; Large Language Models; LLM-Based Reranking
conference: ACM International Conference on Web Search and Data Mining; Feb 22–26, 2026; Boise, USAjournalyear: 2026ccs: Information systems Retrieval models and rankingccs: Computing methodologies Natural language processing
1. Introduction
@@
Figure 1. We extract semantic features, rerank a larger candidate pool with them, and finally refine the shortlist with full text instead of directly reranking with full text.
[…]
Flow diagram with three stages: (i) offline extraction of document-level features; (ii) coarse-grained reranking over a large candidate set using compact feature vectors; (iii) fine-grained reranking of the shortlisted candidates using full-text representation. Arrows indicate the processing order from (i) to (iii).
@@
Scientific retrieval (Lawrence et al., 1999; White et al., 2009) is crucial for scientific knowledge discovery. While current retrievers are effective at retrieving broadly relevant scientific papers, they often struggle to differentiate among documents covering similar topics (Sciavolino et al., 2021; Liu et al., 2021), failing to make fine-grained relevance estimation. Therefore, the reranking stage (Carbonell and Goldstein-Stewart, 1998; Kurland and Lee, 2005) is particularly important (Gao et al., 2021b), as it refines the first-stage retrieval to better distinguish between closely ranked documents.
@@
Recently, Large Language Models (LLMs) (Grattafiori et al., 2024; OpenAI et al., 2024) have significantly advanced the field of document reranking, particularly through their application in listwise reranking (Sun et al., 2023b; Ma et al., 2023). In this setting, LLMs jointly assess a set of retrieved candidates within their context window and generate a reordered ranking based on their relevance. Previous works (Pradeep et al., 2023a; Gangi Reddy et al., 2024; Liu et al., 2024b; Ren et al., 2024) show that LLM-based listwise rerankers outperform prior embedding-based approaches (Nogueira et al., 2019, 2020) on multiple evaluation datasets (Thakur et al., 2021).
@@
The standard practice in LLM listwise reranking (Sun et al., 2023b; Pradeep et al., 2023a, b; Gangi Reddy et al., 2024; Liu et al., 2024b) is to input the full text of each candidate document into the context window. In scientific retrieval, however, this approach faces key limitations. In the scientific domain, first-stage retrievers often show limited performance (Kim et al., 2023; Kang et al., 2024a, b), so truly relevant documents may not rank high enough in the first-stage retrieval. Moreover, due to the substantial token overhead of full text representation and finite context length, rerankers can only operate on a limited number (typically 20 per prompt) (Sun et al., 2023b; Ma et al., 2023) of candidates. Consequently, when the first-stage retrieval is suboptimal, many relevant documents are excluded from reranking consideration, which inherently constrains overall performance (Reddy et al., 2023).
A straightforward solution is to expand the context window of LLM rerankers to handle more candidates. This can be done using long-context LMs (Huang et al., 2023; Pawar et al., 2024) or test-time scaling strategies (Snell et al., 2024; Xia et al., 2025), such as multi-round reranking (Sun et al., 2023b; Ma et al., 2023). However, these methods have notable drawbacks: long-context LMs require additional training (Liu et al., 2024b), suffer from “loss in the middle” problem (Liu et al., 2024a; Hsieh et al., 2024), and incur high computational costs (Liu et al., 2025b); sliding windows similarly increase inference latency and overhead (Sun et al., 2023b; Ma et al., 2023). To better address the limitations, we explore compressing each document’s representation by abstracting high-level features (e.g., categories, keywords), enabling more documents to fit within a fixed context window. This achieves a similar effect to window expansion but with greater efficiency.
@@
We propose CoRank, a zero-shot, model-agnostic reranking framework for science retrieval as exhibited in Figure 1. CoRank consists of three stages: (i) Offline Preprocessing: extract high-level semantic features like categories and keywords from unstructured scientific documents; (ii) Coarse-grained Reranking: use these compact representations for an initial ranking and select a subset of top candidates; (iii) Fine-grained Reranking: rerank the top candidates using full scientific documents to recover the potentially missing details during information extraction. This integrated reranking framework aims to address the suboptimal first-stage retrieval. Compact representations allow more documents to fit within the context window, improving candidate set coverage, while the final fine-grained ranking ensures a more accurate ordering.
We comprehensively evaluate CoRank on five diverse scientific retrieval datasets spanning computer science (Ajith et al., 2024; Mysore et al., 2021), biomedical (Boteva et al., 2016), and clinical (Wadden et al., 2020; Voorhees et al., 2021) domains. Our experiments cover a diversity of LLM backbones, including both proprietary (OpenAI et al., 2024; Google DeepMind, 2025) and open-source (Qwen et al., 2024; Jiang et al., 2023; Grattafiori et al., 2024) models. Results show that CoRank consistently outperforms existing LLM-based zero-shot reranking baselines, achieving an average absolute improvement of +4.9+4.9 nDCG@10. Notably, our method delivers strong gains even in challenging settings with suboptimal first-stage retrieval, while also maintaining high efficiency by reducing token consumption compared to the widely adopted sliding window strategy (Sun et al., 2023b). Extensive ablation studies and further analyses validate the contribution of each component and demonstrate the robustness of our method.
@@
    We reveal the limitations of LLM listwise reranking in scientific retrieval.
@@
    We explore semantic features for document representation in reranking and propose CoRank, a novel zero-shot reranking framework.
@@
    We show that CoRank significantly improves reranking performance of LLM-based listwise approach in scientific retrieval.
These findings underscore how semantic information extraction and retrieval interact, with pre-analysis of document features enabling more effective reranking in scientific search.
2. Related Work
2.1. Classical Document Reranking
Document reranking (Karpukhin et al., 2020; Ren et al., 2021, 2023) is a critical component in information retrieval (IR)  (Baeza-Yates et al., 1999; Singhal et al., 2001), especially when high retrieval precision is required and the first-stage retriever alone is insufficient. Early reranking methods were lexical and probabilistic  (Salton et al., 1975; Robertson and Jones, 1976). Like early sparse retrieval methods (Baeza-Yates et al., 1999; Robertson et al., 2009), they relied on term overlap between the query and document to adjust relevance scores. However, unlike initial retrieval, rerankers revisit a smaller set of candidate documents returned by the first stage and apply more sophisticated weighting or adjustment techniques to refine the ranking. Nevertheless, these approaches were limited by their reliance on exact matching and failed to capture in-depth semantics (Karpukhin et al., 2020; Gao et al., 2021a).
With advances in deep learning, neural reranking methods (Guo et al., 2019; Trabelsi et al., 2021), particularly those based on learning-to-rank (Cao et al., 2007; Liu et al., 2009) frameworks, became prevalent. These models overcome previous limitations, capturing semantic similarity beyond surface-level term matches. The emergence of pre-trained language models (PLMs) (Vaswani et al., 2017; Devlin et al., 2019; Reimers and Gurevych, 2019) further transformed document reranking (Nogueira and Lin, 2019; Yates et al., 2021). Cross-encoder-based rerankers (Litschko et al., 2022), in particular, have demonstrated strong performance by modeling query-document relevance scores and optimizing pointwise, pairwise, or listwise loss (Zhuang et al., 2023). Unlike bi-encoder architectures used in dense retrieval (Karpukhin et al., 2020), cross-encoders jointly encode both the query and each candidate document, allowing the model to directly capture rich dependencies between queries and documents from first-stage retrieval.
@@ 2.1. Classical Document Reranking
Figure 2. Overview of our feature extraction pipeline: from unstructured documents, we apply zero-shot LLM information extraction to obtain document features including categories, sections, pseudo queries, and keywords, which are then combined into compact representations.
[…]
Flow diagram of the document representation pipeline. Left: an unstructured scientific document enters the system. Middle: a zero-shot LLM information extraction step produces four feature sets—categories (three-level path), section titles, pseudo queries, and keywords. An adaptive selection module keeps the most relevant items (e.g., one section, one pseudo query, and top-k keywords). Right: the selected items are concatenated into a compact representation that is passed to the reranker. Arrows indicate the order: document → extraction → selection → compact representation → reranking.
2.2. LLM-Based Listwise Document Reranking
Recently, large language models (LLMs) (Touvron et al., 2023; OpenAI et al., 2023) have revolutionized the field of NLP. With strong general knowledge and instruction-following (Ouyang et al., 2022) capabilities, LLMs offer a new paradigm for listwise document reranking. A classic proposal (Sun et al., 2023b) involves placing the query and a list of candidate documents into the context window and prompting the LLM to generate a reordered ranking. This zero-shot strategy has been shown to outperform the traditional reranking baselines (Sun et al., 2023b; Ma et al., 2023).
After the introduction of LLM-based listwise reranking, research has rapidly extended in several directions. For model development, works such as RankVicuna (Pradeep et al., 2023a) and RankZephyr (Pradeep et al., 2023b) pioneered open-source reranking LLMs, while RankMistral (Liu et al., 2024b) leveraged long-context training to adapt LLMs for larger document lists in a single input. For context-length adaptation, Sliding window strategies (Sun et al., 2023b; Ma et al., 2023) partition candidate lists into overlapping chunks for broader coverage, addressing the limited context windows of LLMs. For inference efficiency, FIRST (Gangi Reddy et al., 2024) improves the speed of listwise reranking by restricting the process to first-token decoding. Liu et al. (2025a) and Li et al. (2024a) use passage embeddings and chunked text blocks to represent documents reducing the token cost per document. From the perspective of self-consistency, ScaLR (Ren et al., 2024) introduces self-calibration to improve consistency across windows. Meanwhile, Tang et al. (2024) enhance reranking quality through permutation consistency.
Since our method operates on document representations in a plug-and-play fashion, it is orthogonal to most previous baselines and can be applied jointly with them (we illustrate this with the sliding window strategy in Section 5). Notably, two prior studies (Liu et al., 2025a; Li et al., 2024a) have also identified limitations of full-document representations, but their approaches are fundamentally different from ours: one (Liu et al., 2025a) adopts non-natural language embeddings, while the other (Li et al., 2024a) merely relies on document chunking. Furthermore, their analyses focus on the general domain, whereas ours is conducted in the scientific domain.
Overall, our work highlights the limitations of full-document representations for listwise reranking in scientific retrieval. And we leverage IE-based features as document representations for reranking. This makes our contribution novel and significant within the reranking literature.
Figure 3. Token efficiency and performance comparison across different document representations. (a) Per-document token length distribution. (b) nDCG@10 scores for different numbers of documents in the context window of single LLM input (c) Context token overhead to reach equal reranking performance.
[…]
Composite figure with three plots. (a) Violin/box plots show token counts per document: full text averages around 200 tokens, with a long tail past 5,000; four feature-based forms (Form 1–4) are far shorter, roughly 10–50 tokens each. (b) Line chart of nDCG@10 as candidate-set size rises from 10 to 200. Full text reranking leads up to  80 documents, but exceeds the 32k-token limit thereafter; Form 4 continues improving and overtakes full text beyond 100 candidates. (c) Bar chart of total context tokens required to reach 70%, 85%, 100%, 115%, and 130% of the 20-doc full text baseline. Full text needs ≈3,500 tokens for 100% performance, whereas Form 4 reaches the same level with ≈1,000 tokens and achieves 130% performance with fewer tokens than full text at 100%. The visual message: feature-based representations are dramatically more token-efficient, allowing more documents per prompt and equal or better reranking quality.
3. Preliminary Analysis
@@
In this section, we first define the task of listwise reranking, then show the limitations of current full-document-based methods and explore compact and informative document representation options for reranking in the scientific domain.
3.1. Problem Definition
@@
Reranking Input. Given a query $q$ and a large corpus $\mathcal{P}_{n}$ of $n$ scientific papers, a first-stage retriever selects a ranked list of $m$ candidate documents $\mathcal{C}=\{p_{1},p_{2},\dots,p_{m}\}\subset\mathcal{P}_{n}$. Typically we have $m\ll n$ due to the scale of the corpus and the finite capacity of rerankers. This candidate list $\mathcal{C}$ is then passed to a reranking model, which aims to reorder the documents such that more relevant ones are ranked higher.
@@
Reranking Objective. The goal of document reranking is to find a permutation $\sigma:\{1,\dots,m\}\!\to\!\{1,\dots,m\}$ whose application
@@
orders documents in more accurate descending relevance to $q$. The quality is often (Sun et al., 2023b; Gangi Reddy et al., 2024; Liu et al., 2024b) evaluated using top-$k$ metrics such as nDCG, with small $k$ values that stress accuracy at the top of the ranked list.
LLM Listwise Reranking. Large language models can inspect $m$ candidate documents in their context window and generate a permutation that reflects their relevance ordering. Formally, the model acts as follows:
$\sigma_{\text{LLM}}\;=\;\mathrm{LLM}(q,\,\mathcal{C})$
Applying this permutation to the candidate list yields
@@
3.2. Current Limitations
The standard approach in LLM-based listwise reranking (Sun et al., 2023b; Pradeep et al., 2023a, b; Gangi Reddy et al., 2024; Liu et al., 2024b) feeds the full text of each candidate into the model’s context window. While this allows LLMs to capture the complete content of each document, it faces some challenges in the scientific domain.
Suboptimal First-Stage Retrieval. Unlike in general-domain scenarios, first-stage retrievers—whether sparse (Robertson et al., 2009; Nogueira and Lin, 2019) or dense (Izacard et al., 2021; Wang et al., 2022)—struggle to generalize to the scientific domain (Thakur et al., 2021; Bonifacio et al., 2022). This is due to the complexity of long-tail concepts (Kang et al., 2024b, 2025) and the lack of large-scale supervised training data (Bonifacio et al., 2022; Li et al., 2023). Consequently, their retrieval performance is limited, and truly relevant documents are often ranked much lower (Ajith et al., 2024).
Token Consumption. Full text representation also introduces significant token overhead, with individual scientific papers often consuming hundreds or even thousands of tokens (Thakur et al., 2021; Ajith et al., 2024; Mysore et al., 2021). At the same time, LLMs have a limited effective context length (Dai et al., 2019; Xiong et al., 2024) and are known to suffer from issues such as positional bias for long inputs (Liu et al., 2024a; Tian et al., 2025). As a result, full text rerankers are constrained to operate on a small number of candidates, typically 20 documents per input prompt (Sun et al., 2023b; Ma et al., 2023; Pradeep et al., 2023a, b).
@@ 3.2. Current Limitations
3.3. Semantic Features as Alternatives
3.3.1. Semantic Feature Construction
Given the limitations of full text representations in scientific-domain reranking, we explore an alternative approach, which is based on document-level semantic features such as categories and keywords. The intuition is that these features are both significantly more concise and capable of preserving the core semantics. Therefore, more candidate documents can be effectively considered during reranking. Specifically, we examine the following four types of document semantic features:
Category. Categories (Sun et al., 2023a; Zhang et al., 2025) offer a high-level topical overview of scientific documents. Each document is assigned a three-level hierarchical category path in the format {Category} → {Subcategory} → {Subsubcategory}, which provides a broad-to-specific classification on its topic.
Section. Sections (Zhou et al., 2023) consist of multiple subtitle-style strings, each summarizing a major part of the scientific document. They capture the internal structure of the document and serve as mid-level semantic signals that enhance category-level summaries.
Keyword. Keywords (Rose et al., 2010; Lee et al., 2023) are terms or entities that represent fine-grained lexical concepts within a document. They provide the most specific information among different granularities.
Pseudo Query. Pseudo queries (Sachan et al., 2022; Kang et al., 2025) are synthetic user questions based on the content. This feature offers a unique query-aligned perspective of the document, simulating how the document might be retrieved in real-world scenarios.
These four features represent the most common types of document-level information extraction (IE) (Niklaus et al., 2018; Wimalasuriya and Dou, 2010), varying in both granularity and style. All semantic features are extracted using LLM-based information extraction methods (Yuan et al., 2023; Zhang et al., 2023). We adopt zero-shot prompting with instruction-tuned models such as GPT-4.1 (OpenAI, 2024) and the Qwen model family (Qwen et al., 2024). This approach leverages the inherent knowledge of pretrained language models (Min et al., 2023; Li et al., 2024b) to extract information without requiring task-specific training (Yuan et al., 2023), while demonstrating strong generalization across domains and schema variations (Xu et al., 2025).
Among these features, sections, keywords, and pseudo queries have multiple elements for each document. To reduce noise and focus on the most relevant content, we apply an adaptive selection strategy during inference time. We compute dense embedding similarities between the query and each extracted element and retain only the most relevant: 5 keywords (from 30), 1 pseudo query (from 20), and 1 section (from 3). This improves content relevance while minimizing token overhead.